OpenStudy (anonymous):

I know I have to do the quadratic equation to unpack this: but I'm having trouble. 14x^2-10x-4356=0

OpenStudy (anonymous):

use this formula 

OpenStudy (anonymous):

I got 18, I think that is correct. Thanks

OpenStudy (academicgurusinc):

Actually you can solve this by factoring which is way easier. So you can take out a "2" and will get 2(x-18)(7x+121)=0 Solve for x-18=0 and you'll get x=18 as one of the answers. Solve for 7x+121=0 and you you'll get 7x=-121 and therefore x=-121/7 as your other answer. I hope that helps. If you found this helpful, I encourage you to subscribe to our youtube channel, to stay current with all of our new videos: https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg You may also want to check out our Math Challenge contest, for a chance to win a $50 Visa gift card: https://www.youtube.com/watch?v=2Dbu-R_Hj7E Regards, Academic Gurus Inc.

OpenStudy (anonymous):

Ok, but how come you can take out a 2, what happens to that 2

OpenStudy (academicgurusinc):

You simply do not use it. You only use what you have inside your parenthesis.

OpenStudy (mathstudent55):

Factoring: \(14x^2-10x-4356=0\) \(2(7x^2-5x-2178 )=0\) Divide both sides by 2. \(7x^2 - 5x - 2178 = 0\) \(\dfrac{(7x - 126)(7x + 121)}{7} = 0\) \((x - 18)(7x - 121) =0\) \(x = 18\) or \(7x = 121\) \(x = 18 \) or \(x = -\dfrac{121}{7} \) Quadratic formula: \(14x^2-10x-4356=0\) Factor out a 2 and divide both sides by 2 like above. \(7x^2 - 5x - 2178 = 0\) \(x = \dfrac{-(-5) \pm \sqrt{(-5)^2 - 4(7)(-2178)}}{2(7)} \) \(x = \dfrac{5 \pm \sqrt{25 +60984}}{14} \) \(x = \dfrac{5 \pm \sqrt{61009}}{14} \) \(x = \dfrac{5 \pm 247}{14} \) \(x = \dfrac{252}{14}\) or \(x = - \dfrac{242}{14} \) \(x = 18\) or \(x = -\dfrac{121}{7} \)

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Thanks everyone:) It helped a lot.